OpenStudy (anonymous):

find the intercepts for the graph of the equation given. x+7y=0 the x intercept is ?

OpenStudy (anonymous):

@aurorablue01

OpenStudy (anonymous):

What value is y when x is zero? Likewise, what value is x when y is zero

OpenStudy (anonymous):

The x intercept would be where this line intersects the x axis, and that is when y is zero

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

What value did you get?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

All you have to do is let y = 0 and substitute, x + 7(0) = x + 0 = x = 0

OpenStudy (anonymous):

So the point is (0,0)